rani father's age is 3 times old as rani.4 years ago ,he was 4 times older.how old is rani?
Answer by ankor@dixie-net.com(22740)

rani father's age is 3 times old as rani.
f = 3r
:
4 years ago ,he was 4 times older.
f - 4 = 4(r-4)
f - 4 = 4r - 16
:
Replace f with 3r
3r - 4 = 4r - 16
-4 = 16 = 4r - 3r
r = 12
:
Check solution in the statement: (f = 36)
"4 years ago ,he was 4 times older.
36 - 4 = 4(12 - 4)
32 - 4(8)